OpenStudy (saylilbaby):

multiply and simplify if possible. raticle sign 7x(raticle sign x-7 raticle sign 7)

OpenStudy (0487308):

Please write the equation numerically.

OpenStudy (saylilbaby):

\[\sqrt{7x}(\sqrt{x}-7\sqrt{7}\] @0487308

OpenStudy (0487308):

Thanks, but I don't know how to solve that ::(

OpenStudy (helpsos):

what do think the first step would be to solving? Do you know the FOIL method?